OpenStudy (anonymous):

Please use substitution to solve this system of equations. 2x-3y=-1 y=x-1 A. (-5,1) B.(4,3) C.(-2,-3) D.(5,-3)

OpenStudy (anonymous):

show ur work

OpenStudy (anonymous):

I did.. yeah.. It just.. it was too much...

OpenStudy (anonymous):

I got B, is that right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let's take the second equation: y=x-1 then plug this x value on the first equation: 2x-3(x-1) = -1 2x - 3x + 3 = -1 -x = -4 x = 4 now plug it back on second equation: y = 4-1 y = 3 SO (4, 3)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

i checked it by plugging the value back into the equations

OpenStudy (anonymous):

its a useful tip

OpenStudy (anonymous):

Thank you.

OpenStudy (texaschic101):

2x - 3y = -1 y = x - 1 now sub x - 1 in for y in the other equation 2x - 3(x-1) = -1 -- distribute 2x - 3x + 3 = -1 -- subtract 3 2x - 3x = -1 - 3 -- combine like terms -x = -4 x = 4 now sub 4 in for x y = x - 1 y = 4 - 1 y = 3 yep....answer is B (4,3)

OpenStudy (anonymous):

Please solve this system of equations. x=y-6 3x+4y=10 A.(0,6) B.(-2, 4) C.(3,9) D.(1,9) So This one, The answer is B aswell?